OpenStudy (anonymous):

An object with a mass of 7.5g raises the water level in the cylinder from 25.1mL to 30.1mL. What is the density of the object

OpenStudy (taramgrant0543664):

Density=mass divided by the final volume

OpenStudy (anonymous):

@Alexa_Noelle33 What do you think the volume of the object is ?

OpenStudy (jhannybean):

\[\sf \rho = \frac{m}{v}~~~~~~~~~~~~~~\rho = density \]\[\sf \rho = \frac{m_{obj}}{\Delta v}~~~~~~~~~\Delta v = change ~in~volume~of~water\]\[\sf \rho = \frac{7.5~g}{(30.1~ml ~-~25.1~ml)}~ =~? ~~\frac{g}{mL}\]

OpenStudy (anonymous):

The initial volume was 25.1mL and then the object of the mass which is 7.5g raised the final volume to 30.1mL. Looking for the density. Density =m/v. Mass given=7.5g. Water level before object submerged =25.1mL. Water level after submerged =30.2mL. Need density (g/mL). 30.1mL-25.1mL=5mL. 7.5g/5mL= 1.5g/mL.

OpenStudy (jhannybean):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @mbamelun The initial volume was 25.1mL and then the object of the mass which is 7.5g raised the final volume to 30.1mL. Looking for the density. Density =m/v. Mass given=7.5g. Water level before object submerged =25.1mL. \(\sf \color{red}{Water ~level ~after ~submerged =30.2mL}\). Need density (g/mL). 30.1mL-25.1mL=5mL. 7.5g/5mL= 1.5g/mL. \(\color{#0cbb34}{\text{End of Quote}}\) Typo there. :P